# Training a neural netwok for data reduction and better generalization

### v2

@@ Abstract
The motivation for sparse learners is to compress the inputs (features) by selecting only the ones needed for good generalization. Linear models with LASSO-type regularization achieve this by setting the weights of irrelevant features to zero, effectively identifying and ignoring them.
In artificial neural networks, this selective focus can be achieved by pruning the input layer. Given a cost function enhanced with a sparsity-promoting penalty, our proposal selects a regularization term $\lambda$ (without the use of cross-validation or a validation set) that creates a local minimum in the cost function at the origin where no features are selected. This local minimum acts as a baseline, meaning that if there is no strong enough signal to justify a feature inclusion, the local minimum remains at zero with a high prescribed probability. The method is flexible, applying to complex models ranging from shallow to deep artificial neural networks and supporting various cost functions and sparsity-promoting penalties.
We empirically show a remarkable phase transition in the probability of retrieving the relevant features, as well as good generalization thanks to the choice of $\lambda$, the non-convex penalty and the optimization scheme developed. This approach can be seen as a form of compressed sensing for complex models, allowing us to distill high-dimensional data into a compact, interpretable subset of meaningful features.
@@ 1.1 Compressed sensing
where $\|{\bm{\theta}}\|_{0}=:s$ counts the number $s$ of non-zero entries of ${\bm{\theta}}$. Solving this discrete optimization problem is computationally intractable. Instead Basis Pursuit (Chen et al., 1999) calculates $\hat{\bm{\theta}}^{\rm BP}$ solution to the continuous optimization problem:
@@ 1.3 Noisy compressed sensing
In practice, the vector ${\bf y}$ is known with random errors. The standard linear model assumes ${\bf y}=c{\bf 1}+X{\bm{\theta}}^{*}+{\bf e}$, where $c$ is the intercept and ${\bf e}$ is a realization from ${\rm N}({\bf 0},\sigma^{2}I_{n})$. Basis Pursuit Denoising extends Basis Pursuit (1) to the noisy scenario, also known as the LASSO (Tibshirani, 1996), by solving
@@
Other generalizations of Basis Pursuit include square root-LASSO (Belloni et al., 2011) for $q=1$, group LASSO for $\nu=2$ (Yuan and Lin, 2006) and LASSO-zero (Descloux and Sardy, 2021). Under some regularity conditions, such approaches retrieve ${\cal S}^{*}$ despite the noise and the high-dimensionality (Bühlmann and van de Geer, 2011). Yet, Su et al. (2015) proved that false discoveries occur early on the LASSO path, causing the phase transition to drop to zero early. To push the phase transition further than with LASSO, Chartrand (2007) and Chartrand and Yin (2008) studied and provided algorithms that replace the $\ell_{1}$ sparsity inducing penalty by the continuous, but non-convex, $\ell_{\nu}$ penalty with $\nu<1$. The idea it to get closer to the $\ell_{0}$ penalty than with the $\ell_{1}$ penalty by letting $\nu$ get close to zero but strictly positive to deal with a continuous penalty.
@@ 1.4 Noisy compressed sensing beyond linearity
Definition 1 (Feature selection compatible ANN).
@@ 2.1 Sparsity inducing penalty
where $\mu_{({\bm{\theta}}^{(1)},{\bm{\theta}}^{(2)})}$ is a learner, ${\cal L}_{n}$ is some goodness-of-fit measure (e.g., negative log-likelihood function, cross-entropy) also called the loss function and $P$ is a sparsity inducing penalty (Bach et al., 2011). The complete expression is referred to as the cost function. Observe that we do not penalize all coefficients since ${\bm{\theta}}^{(2)}$ is not an argument of $P$; this allows for instance the intercept in linear regression to not be penalized (that is, ${\bm{\theta}}^{(2)}=c$). We now give a new definition of a sparsity inducing penalty and introduce the concept of a local zero-thresholding function.
Definition 2 (Sparsity inducing penalty and local zero-thresholding function).
@@ 2.2 Selection of $\lambda$
Selecting the parameter $\lambda$ is crucial for good recovery of the parameters. The typical — though suboptimal — approach to setting the penalty parameter involves splitting the dataset for cross-validation. But, this method tends to be conservative, often detecting too many false features, and is computationally expensive. Instead, the quantile universal threshold (QUT) (Giacobino et al., 2017a) selects a single $\lambda$ for a learner solution to (9) with a sparsity inducing penalty. The idea is to select $\lambda$ at the detection edge between noise and signal by calibrating $\lambda$ correctly under the pure noise model, that is, when the true parameters are ${\bm{\theta}}^{(1)*}={\bf 0}$ and the underlying association $\mu_{({\bm{\theta}})}({\bf x})=c$ is constant. QUT is formally defined as follows.
Definition 3 (Quantile Universel Treshold).
@@
In other words, choosing $\lambda=\lambda_{\alpha}^{\rm QUT}(X)$ in (9) creates a local minimum at $\hat{\bm{\theta}}^{(1)}={\bf 0}$ with probability $1-\alpha$ when the data come from the null model ${\bm{\theta}}^{(1)*}={\bf 0}$. For certain linear models, this way of selecting $\lambda$ by $\lambda_{\alpha}^{\rm QUT}(X)$ allows to retrieve the sign of ${\bm{\theta}}^{(1)*}$ also under sparse alternatives ${\bm{\theta}}^{(1)*}\neq{\bf 0}$ (Donoho and Johnstone, 1994; Donoho et al., 1995; Bühlmann and van de Geer, 2011; Giacobino et al., 2017a).
@@ 2.3 Compatible cost functions
The pair loss and penalty $({\cal L}_{n},P)$ in the cost function (9) must be chosen with care. Not all pairs have a statistic $\Lambda=\lambda_{0}({\bf Y},X)$ that is pivotal—meaning its distribution does not depend on unkown parameters. For instance, for the linear model $\min_{c\in{\mathbb{R}},{\bm{\theta}}\in{\mathbb{R}}^{p}}\|{\bf y}-c{\bf 1}-X{\bm{\theta}}\|_{2}^{q}+\lambda\|{\bm{\theta}}\|_{1}$ with ${\bf y}=c{\bf 1}+X{\bm{\theta}}^{*}+{\bf e}$ and ${\bf e}$ is a realization from ${\rm N}({\bf 0},\sigma^{2}I_{n})$, the corresponding $\Lambda$ is pivotal for $q=1$, but not pivotal for $q=2$. Indeed, for $q=1$, $\lambda_{0}({\bf y},X)=\|X^{\rm T}({\bf y}-\bar{\bf y}{\bf 1})\|_{\infty}/\|({\bf y}-\bar{\bf y}{\bf 1})\|_{2}$ and the ratio makes the statistic $\Lambda$ independent of the standard deviation $\sigma$ of the noise since both numerator and denominator are proportional to $\sigma$. For $q=2$, $\lambda_{0}({\bf y},X)=2\|X^{\rm T}({\bf y}-\bar{\bf y}{\bf 1})\|_{\infty}$ the statistic $\Lambda$ can be used solely provided $\sigma$ is known or can be consistently estimated, which is rare in practice.
@@
For the linear model, the convex $\ell_{1}$ penalty is QUT-compatible for $q=1$ and $q=2$ but the $\ell_{2}$ penalty is not since Sardy and Ma (2024) showed that $\lambda_{0}^{\rm local}({\bf y},X)=1$ when $X$ is orthonormal. Furthermore, the nonconvex $\ell_{\nu}$ Subbotin penalty $P({\bm{\theta}})=\sum_{j}|\theta_{j}|^{\nu}$ with $\nu\in(0,1)$ (Sardy, 2009; Woodworth and Chartrand, 2016), induce sparsity but $P$ is not QUT-compatible. Indeed, since the derivative of $|\theta|^{\nu}$ is +$\infty$+ ∞ at $\theta=0^{+}$ (hence creating a local minimum at the origin) the corresponding local zero-thresholding function $\lambda_{0}^{\rm local}({\bf y},X)=0$ is constant.
@@ 2.4 A class of penalties $P_{\nu}$
For $\nu=1$ it amounts to LASSO’s $\ell_{1}$ convex penalty (albeit a factor $1/2$) and for $\nu\rightarrow 0$ it tends to the $\ell_{0}$ discrete penalty for $|\theta|$ large. The following theorem illustrates in the univariate case with the $\ell_{2}$-loss how the new penalty compares to the hard- and soft-thresholding functions that provide a closed form expression to the $\ell_{0}$- and $\ell_{1}$-penalized least squares, respectively (Donoho and Johnstone, 1994). Considering the univariate case is important because the optimization scheme used to solve (9) in Section (2.5) involves using the univariate solutions iteratively until convergence to a stationary point.
Theorem 1 (Univariate thresholding).
@@
Figure (1) plots the thresholding function $\eta_{\lambda}(y;\nu)$ for $(\lambda,\nu)=(4,0.1)$: clearly visible, the corresponding threshold $\varphi(4,0.1)=2.38$ and jump $\kappa(4,0.1)=1.70$ solution to (13). The parameter $\nu=0.1$ being close to zero, one sees that $\eta_{\lambda}(y;\nu)$ approximates the hard thresholding function $\eta_{\varphi_{0}}(y;{\rm hard})=y\cdot 1_{\{|y|>\varphi_{0}\}}$ with $\varphi_{0}=\varphi(4,0.1)$. With $\nu$ equal to one, $\eta_{\lambda}(y;\nu)$ is the soft thresholding function $\eta_{\varphi_{1}}(y;{\rm soft})=\operatorname*{sign}(y)\cdot\max(|y|-\varphi_{1},0)$ with $\varphi_{1}=\lambda/2$ which corresponds to the $\ell_{1}$ convex penalty albeit the factor $1/2$. Since we aim at approaching the hard thresholding function in a continuous way by letting $\nu$ be small, we call the corresponding method harderLASSO. It also bears its name from the fact that the corresponding optimization is harder to solve when $\nu$ gets close to zero.
@@
Theorem 2 (QUT compatibility of penalties $P_{\nu}$).
Consider (9) with a feature selection compatible MLP as described in Definition 1, with a loss ${\cal L}_{n}\in{\cal C}^{2}$ and with $P({\bm{\theta}}^{(1)})=P_{\nu}({\bm{\theta}}^{(1)})$ defined in (10). Then, for any fixed $\nu\in(0,1]$, the optimization problem has a local minimum at $\hat{\bm{\theta}}^{(1)}={\bf 0}$ if and only if $\lambda\geq\lambda_{0}^{\rm local}({\bf y},X)$ for some finite data-dependent value $\lambda_{0}({\bf y},X)$.
The proof of Theorem 2 is given in Appendix B. Ma et al. (2022) provide the closed form expression of the local zero-thresholding function $\lambda_{0}({\bf y},X)$ in regression and classification. Using the square root MSE loss $\mathcal{L}_{n}=\|\mathbf{y}-\mu_{\bm{\theta}}(X)\|_{2}$, the local zero-thresholding function is of the form
@@ 2.5 Optimization
Because of the non-differentiable penalty term of the cost function (9) to fit harderLASSO to the data, a gradient descent-based algorithm like Adam cannot be employed as such, but must be combined with an algorithm specifically designed for dealing with a sparsity inducing penalty like ISTA (Beck and Teboulle, 2009). Applying an ISTA step to minimize an optimization problem of the form $f(\theta)+g(\theta)$ requires solving
@@
at each iteration to move from the current iterate $\bm{\theta}_{(k)}$ to $\bm{\theta}_{(k+1)}$, where $\delta_{(k)}$ is a step size (similar to a learning rate). In our case, $g$ plays the role of the nondifferentiable sparsity inducing penalty $g(\bm{\theta})=\lambda P_{\nu}\left(\bm{\theta}^{(1)}\right)$ given in (10). The multivariate optimization (14) is separable and amounts to solving $\gamma=\sum_{k=1}^{l}p_{k+1}(p_{k}+1)$ univariate problems, either least squares problems for $\bm{\theta}^{(2)}$ or the univariate problems of Theorem 1 for $\bm{\theta}^{(1)}$.
@@
During the training with intermediate $(\lambda_{k},\nu_{k})$ pairs, we employ the Adam optimizer with an initial learning rate of $0.01$, while keeping the remaining optimizer parameters at their default settings. During training, the loss is monitored, and the learning rate is reduced by a factor of $0.5$ if an increase in loss is detected. In the final training phase, only the penalized parameters are updated. Due to the differing characteristics of the weights in the first layer and the biases, a block coordinate relaxation approach is adopted (Tseng, 2001; Sardy et al., 2000). Specifically, blocks are updated successively: first, the biases $(\mathbf{b}_{1},\ldots,\mathbf{b}_{l-1})$ are updated, followed by the first layer weights $W_{1}$, with each block solved using ISTA. A line search determines the step size for each block, after which momentum is applied to the updates. Each training phase is concluded when the relative improvement in the cost falls below a small threshold.

### v3

# Training a neural network for data interpretation and better generalization: towards intelligent artificial intelligence

@@ Abstract
At the time of environmental concerns about artificial intelligence, in particular its need for greedy storage and computation, sparsity inducing neural networks offer a promising path towards frugality and solution for less waste.
Sparse learners compress the inputs (features) by selecting only the ones needed for good generalization. A human scientist can then give an intelligent interpretation to the few selected features. If genes are the inputs and cancer type is the output, then the selected genes give the cancerologist clues on what genes have an effect on certain cancers. LASSO-type regularization leads to good input selection for linear associations, but few attempts have been made for nonlinear associations modeled as an artificial neural network. A stringent but efficient way of testing whether a feature selection method works is to check if a phase transition occurs in the probability of retrieving the relevant features, as observed and mathematically studied for linear models. Our method achieves just so for artificial neural networks, and, on real data, it has the best compromise between number of selected features and generalization performance.
Our method is flexible, applying to complex models ranging from shallow to deep artificial neural networks and supporting various cost functions and sparsity-promoting penalties. It does not rely on cross-validation or on a validation set to select its single regularization parameter making it user-friendly. Our approach can be seen as a form of compressed sensing for complex models, allowing to distill high-dimensional data into a compact, interpretable subset of meaningful features, just the opposite of a black box.
A python package is available at https://github.com/VcMaxouuu/AnnHarderLasso containing all the simulations and ready-to-use models.
@@ 1.1 Compressed sensing
where $\|{\bm{\theta}}\|_{0}=:s$ counts the number $s$ of non-zero entries of ${\bm{\theta}}$. Solving this discrete optimization problem is computationally intractable. Instead Basis Pursuit Chen et al. (1999) calculates $\hat{\bm{\theta}}^{\rm BP}$ solution to the continuous optimization problem:
@@ 1.3 Noisy compressed sensing
In practice, the vector ${\bf y}$ is known with random errors. The standard linear model assumes ${\bf y}=c{\bf 1}+X{\bm{\theta}}^{*}+{\bf e}$, where $c$ is the intercept and ${\bf e}$ is a realization from ${\rm N}({\bf 0},\sigma^{2}I_{n})$. Basis Pursuit Denoising extends Basis Pursuit (1) to the noisy scenario, also known as the LASSO Tibshirani (1996), by solving
@@
Other generalizations of Basis Pursuit include square root-LASSO Belloni et al. (2011) for $q=1$, group LASSO for $\nu=2$ Yuan and Lin (2006) and LASSO-zero Descloux and Sardy (2021). Under some regularity conditions, such approaches retrieve ${\cal S}^{*}$ despite the noise and the high-dimensionality Bühlmann and van de Geer (2011). Yet, Su et al. (2015) proved that false discoveries occur early on the LASSO path, causing the phase transition to drop to zero early. To push the phase transition further than with LASSO, Chartrand (2007) and Chartrand and Yin (2008) studied and provided algorithms that replace the $\ell_{1}$ sparsity inducing penalty by the continuous, but non-convex, $\ell_{\nu}$ penalty with $\nu<1$. The idea it to get closer to the $\ell_{0}$ penalty than with the $\ell_{1}$ penalty by letting $\nu$ get close to zero but strictly positive to deal with a continuous penalty.
@@ 1.4 Noisy compressed sensing beyond linearity
Definition 1.1 (Feature selection compatible ANN).
@@ 2.1 Sparsity inducing penalty
where $\mu_{({\bm{\theta}}^{(1)},{\bm{\theta}}^{(2)})}$ is a learner, ${\cal L}_{n}$ is some goodness-of-fit measure (e.g., negative log-likelihood function, cross-entropy) also called the loss function and $P$ is a sparsity inducing penalty Bach et al. (2011). The complete expression is referred to as the cost function. Observe that we do not penalize all coefficients since ${\bm{\theta}}^{(2)}$ is not an argument of $P$; this allows for instance the intercept in linear regression to not be penalized (that is, ${\bm{\theta}}^{(2)}=c$). We now give a new definition of a sparsity inducing penalty and introduce the concept of a local zero-thresholding function.
Definition 2.1 (Sparsity inducing penalty and local zero-thresholding function).
@@ 2.2 Selection of $\lambda$
Selecting the parameter $\lambda$ is crucial for good recovery of the parameters. The typical-though suboptimal-approach to setting the penalty parameter involves splitting the dataset for cross-validation. But, this method tends to be conservative, often detecting too many false features, and is computationally expensive. Instead, the quantile universal threshold (QUT) Giacobino et al. (2017a) selects a single $\lambda$ for a learner solution to (9) with a sparsity inducing penalty. The idea is to select $\lambda$ at the detection edge between noise and signal by calibrating $\lambda$ correctly under the pure noise model, that is, when the true parameters are ${\bm{\theta}}^{(1)*}={\bf 0}$ and the underlying association $\mu_{({\bm{\theta}})}({\bf x})=c$ is constant. QUT is formally defined as follows.
Definition 2.2 (Quantile Universal Threshold).
@@
In other words, choosing $\lambda=\lambda_{\alpha}^{\rm QUT}(X)$ in (9) creates a local minimum at $\hat{\bm{\theta}}^{(1)}={\bf 0}$ with probability $1-\alpha$ when the data come from the null model ${\bm{\theta}}^{(1)*}={\bf 0}$. For certain linear models, this way of selecting $\lambda$ by $\lambda_{\alpha}^{\rm QUT}(X)$ allows to retrieve the sign of ${\bm{\theta}}^{(1)*}$ also under sparse alternatives ${\bm{\theta}}^{(1)*}\neq{\bf 0}$ Donoho and Johnstone (1994); Donoho et al. (1995); Bühlmann and van de Geer (2011); Giacobino et al. (2017a).
@@ 2.3 Compatible cost functions
The pair loss and penalty $({\cal L}_{n},P)$ in the cost function (9) must be chosen with care. Not all pairs have a statistic $\Lambda=\lambda_{0}({\bf Y},X)$ that is pivotal-meaning its distribution does not depend on unkown parameters. For instance, for the linear model $\min_{c\in{\mathbb{R}},{\bm{\theta}}\in{\mathbb{R}}^{p}}\|{\bf y}-c{\bf 1}-X{\bm{\theta}}\|_{2}^{q}+\lambda\|{\bm{\theta}}\|_{1}$ with ${\bf y}=c{\bf 1}+X{\bm{\theta}}^{*}+{\bf e}$ and ${\bf e}$ is a realization from ${\rm N}({\bf 0},\sigma^{2}I_{n})$, the corresponding $\Lambda$ is pivotal for $q=1$, but not pivotal for $q=2$. Indeed, for $q=1$, $\lambda_{0}({\bf y},X)=\|X^{\rm T}({\bf y}-\bar{\bf y}{\bf 1})\|_{\infty}/\|({\bf y}-\bar{\bf y}{\bf 1})\|_{2}$ and the ratio makes the statistic $\Lambda$ independent of the standard deviation $\sigma$ of the noise since both numerator and denominator are proportional to $\sigma$. For $q=2$, $\lambda_{0}({\bf y},X)=2\|X^{\rm T}({\bf y}-\bar{\bf y}{\bf 1})\|_{\infty}$ the statistic $\Lambda$ can be used solely provided $\sigma$ is known or can be consistently estimated, which is rare in practice.
@@
For the linear model, the convex $\ell_{1}$ penalty is QUT-compatible for $q=1$ and $q=2$ but the $\ell_{2}$ penalty is not since Sardy and Ma (2024) showed that $\lambda_{0}^{\rm local}({\bf y},X)=1$ when $X$ is orthonormal. Furthermore, the nonconvex $\ell_{\nu}$ Subbotin penalty $P({\bm{\theta}})=\sum_{j}|\theta_{j}|^{\nu}$ with $\nu\in(0,1)$ Sardy (2009); Woodworth and Chartrand (2016), induce sparsity but $P$ is not QUT-compatible. Indeed, since the derivative of $|\theta|^{\nu}$ is +$\infty$+ ∞ at $\theta=0^{+}$ (hence creating a local minimum at the origin) the corresponding local zero-thresholding function $\lambda_{0}^{\rm local}({\bf y},X)=0$ is constant.
@@ 2.4 A class of penalties $P_{\nu}$
For $\nu=1$ it amounts to LASSO’s $\ell_{1}$ convex penalty (albeit a factor $1/2$) and for $\nu\rightarrow 0$ it tends to the $\ell_{0}$ discrete penalty for $|\theta|$ large. The following theorem illustrates in the univariate case with the $\ell_{2}$-loss how the new penalty compares to the hard- and soft-thresholding functions that provide a closed form expression to the $\ell_{0}$- and $\ell_{1}$-penalized least squares, respectively Donoho and Johnstone (1994). Considering the univariate case is important because the optimization scheme used to solve (9) in Section (2.5) involves using the univariate solutions iteratively until convergence to a stationary point.
Theorem 2.4 (Univariate thresholding).
@@
Figure (1) plots the thresholding function $\eta_{\lambda}(y;\nu)$ for $(\lambda,\nu)=(4,0.1)$: clearly visible, the corresponding threshold $\varphi(4,0.1)=2.38$ and jump $\kappa(4,0.1)=1.70$ solution to (13). The parameter $\nu=0.1$ being close to zero, one sees that $\eta_{\lambda}(y;\nu)$ approximates the hard thresholding function $\eta_{\varphi_{0}}(y;{\rm hard})=y\cdot 1_{\{|y|>\varphi_{0}\}}$ with $\varphi_{0}=\varphi(4,0.1)$. With $\nu$ equals to one, $\eta_{\lambda}(y;\nu)$ is the soft thresholding function $\eta_{\varphi_{1}}(y;{\rm soft})=\operatorname*{sign}(y)\cdot\max(|y|-\varphi_{1},0)$ with $\varphi_{1}=\lambda/2$ which corresponds to the $\ell_{1}$ convex penalty albeit the factor $1/2$. Since we aim at approaching the hard thresholding function in a continuous way by letting $\nu$ be small, we call the corresponding method harderLASSO. It also bears its name from the fact that the corresponding optimization is harder to solve when $\nu$ gets close to zero.
@@
Theorem 2.5 (QUT compatibility of penalties $P_{\nu}$).
Consider (9) with a feature selection compatible MLP as described in Definition 1.1, with a loss ${\cal L}_{n}\in{\cal C}^{2}$ and with $P({\bm{\theta}}^{(1)})=P_{\nu}({\bm{\theta}}^{(1)})$ defined in (10). Then, for any fixed $\nu\in(0,1]$, the optimization problem has a local minimum at $\hat{\bm{\theta}}^{(1)}={\bf 0}$ if and only if $\lambda\geq\lambda_{0}^{\rm local}({\bf y},X)$ for some finite data-dependent value $\lambda_{0}({\bf y},X)$.
The proof of Theorem 2.5 is given in Appendix B. Ma et al. (2022) provide the closed form expression of the local zero-thresholding function $\lambda_{0}({\bf y},X)$ in regression and classification. Using the square root MSE loss $\mathcal{L}_{n}=\|\mathbf{y}-\mu_{\bm{\theta}}(X)\|_{2}$, the local zero-thresholding function is of the form
@@ 2.5 Optimization
Because of the non-differentiable penalty term of the cost function (9) to fit harderLASSO to the data, a gradient descent-based algorithm like Adam cannot be employed as such, but must be combined with an algorithm specifically designed for dealing with a sparsity inducing penalty like ISTA Beck and Teboulle (2009). Applying an ISTA step to minimize an optimization problem of the form $f(\theta)+g(\theta)$ requires solving
@@
at each iteration to move from the current iterate $\bm{\theta}_{(k)}$ to $\bm{\theta}_{(k+1)}$, where $\delta_{(k)}$ is a step size (similar to a learning rate). In our case, $g$ plays the role of the nondifferentiable sparsity inducing penalty $g(\bm{\theta})=\lambda P_{\nu}\left(\bm{\theta}^{(1)}\right)$ given in (10). The multivariate optimization (14) is separable and amounts to solving $\gamma=\sum_{k=1}^{l}p_{k+1}(p_{k}+1)$ univariate problems, either least squares problems for $\bm{\theta}^{(2)}$ or the univariate problems of Theorem 2.4 for $\bm{\theta}^{(1)}$.
@@
During the training with intermediate $(\lambda_{k},\nu_{k})$ pairs, we employ the Adam optimizer with an initial learning rate of $0.01$, while keeping the remaining optimizer parameters at their default settings. During training, the loss is monitored, and the learning rate is reduced by a factor of $0.5$ if an increase in loss is detected. In the final training phase, only the penalized parameters are updated. Due to the differing characteristics of the weights in the first layer and the biases, a block coordinate relaxation approach is adopted Tseng (2001); Sardy et al. (2000). Specifically, blocks are updated successively: first, the biases $(\mathbf{b}_{1},\ldots,\mathbf{b}_{l-1})$ are updated, followed by the first layer weights $W_{1}$, with each block solved using ISTA. A line search determines the step size for each block, after which momentum is applied to the updates. Each training phase is concluded when the relative improvement in the cost falls below a small threshold.